INAF queue observing, 2017 Feb 16-17

Observer: R. Carini, F. Cusano
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: D. Thompson on call
Instrument:

Summary

At the sunset some clouds on the horizon

Night duration (12\xB0 el twilight) 02:30am-05:45pm UT) = 10 h 15m

Observing time: 1 hrs 0 m (10%)

Weather time loss: 8 hrs 55m (87%)

technical time loss 20 m (3%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Phot. Standard S98 blue: S g r; red: V I r    
standards star gd71 dual grating    
Antoniucci IPTF15AFQ dual grating 0.5 h only spectra (mods1+mods2)

Details

01:56 photometric standard
  • dohybrid
  • S98_Sgr_VIr.ob
  • seeing 2.0'', clouds are coming

02:17 D'Avanzo program, target GRB 160408A
  • dofpia, X2
  • we stop the dofpia because the sky is completely covered.

02:20 we switch to MODS

02:50 the seeing is so bad (> 2.0'') that is impossible the collimate

03:27 finally the collimate has done. We send the acquisition for the standard star
  • acqBinoMODS gd71.acq mods1r.20170217.0002-0003 mods2r.20170217.0002-0004
  • seeing about 2.2'''
  • 03:30 problems with modsAlign in MODS2: "An internal error has been detected: invalid command name"
  • 03:46 we do the aligment by hand
  • 03:54 execBinoMODS gd71.obs
  • seeing about 2.5'' mods1r.0004-0005, mods1b.0002-0003 mods2r.0005-0006, mod2b.0002-0003

04:00 slew to Antoniucci Backup program, target IPTF15AFQ
  • send the acquisition
  • lost guide star
  • send the acquisition again: acqBinoMODS iptf15afq.acq mods1r.0006-0008 mods2r.0007-0009
  • 04:15 same error during the modsAlign also with MODS1
  • 04:22 the SA send the modsAling from the control room, it works. the SA thinks that the ds9 opened with LBC was in conflict with ds9 opened with MODS
  • 04:27 execBinoMODS iptf15afq.obs mods1r.0004-0005, mods1b.0002-0003 mods2r.0005-0006, mod2b.0002-0003
  • seeing about 2.0''-3.0''

04:54 the TO closes the dome due to the clouds

06:05 switch to LUCI1

06:29 we begin Marziani program, target C. the TO send the preset.
  • the seeing is horrible, there are problems with the collimation
  • 06:56 clouds come again rapidly
  • 07:00 the TO closes again the dome

-- %USERSIG{RobertaCarini - 2017-02-17}%

Comments

 
Topic revision: r9 - 24 Jan 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback